OpenStudy (anonymous):

2 y^3-15y=y^2 value of this by factoring? work.

OpenStudy (jdoe0001):

\(\bf 2 y^3-15y=y^2 \ \ \ ?\)

OpenStudy (akashdeepdeb):

Yes I think that's what it is! :D

OpenStudy (anonymous):

How about x^2+2x-8=0